# SeaMo: A Multi-Seasonal and Multimodal Remote Sensing Foundation Model

### v1

@@ 3.1.2 Mask operation
Next, a masking operation is conducted on the patches to distinguish between unmasked and masked patches. This is represented as $[\bm{{O}}^{t}_{visible},\bm{{O}}^{t}_{masked}]=\mathbb{M}\odot\bm{{E}}_{O}^{t}$ and $[\bm{{R}}^{t}_{visible},\bm{{R}}^{t}_{masked}]=\mathbb{M}\odot\bm{{E}}_{R}^{t}$, where $\mathbb{M}\in\{0,1\}^{L\times D}$ is a binary mask that indicates the patches to be masked. Notably, we employ a consistent masking strategy for images from different sensors captured at the same time and a random masking strategy for images from the same sensor but captured at different times. This approach enhances the reconstruction challenge, encouraging the model to learn more robust representations.
3.1.3 Encoder
We concatenate the visible tokens from two modalities and input them into the encoder $\Phi$, denoted as $\{\bm{F}^{t}_{O},\bm{F}^{t}_{R}\}=\Phi(\text{concat}[\bm{{O}}^{t}_{visible},\bm{{R}}^{t}_{visible}])$. This encoder utilizes plain vision transformer blocks [37]. Following the random masking process, $\bm{{O}}^{t}_{visible}$ and $\bm{{R}}^{t}_{visible}$ represent tokens that vary not only in their modality but also in spatial attributes. Consequently, the encoder’s self-attention mechanism is capable of extracting features and building the relationships of all visible tokens between multimodal and spatial dimensions.
@@ 3.1.3 Encoder
Figure 3: An illustration of the Temporal-Multimodal (TM) block. In this block, data from each modality not only engages in data fusion interactions at the current season but also influences the fusion process in subsequent seasons.
3.1.4 Temporal-Multimodal fusion block
The output from the encoder generates multimodal features at distinct time points, represented as $\{\bm{F}^{t}_{O},\bm{F}^{t}_{R}\},t=1,2,...,T$. Although seasonal images are processed through the encoder, the neural network cannot currently explicitly learn time-invariant representations, leaving temporal features underutilized. To address this, we introduce the Temporal-Multimodal (TM) fusion block (see Figure 3), inspired by SiamMAE [38], which is designed to explicitly learn temporal correspondences in videos. In this block, each modality data participates not only in the interaction of data fusion at the current time point but also contributes to the fusion at the subsequent time point. Algorithm 1 provides the pseudocode for the Temporal-Multimodal (TM) fusion block. Since all images are to be reconstructed, the visible tokens of each image act as the query. Tokens from other modalities of the same season, combined with tokens from the previous season, serve as the key and value. Cross-attention is then employed to facilitate interactions across temporal and modal dimensions.
3.1.5 Decoder
The output from the TM block consists of features from distinct modalities at each time step, denoted as $\{\bm{{H}}^{t}_{O},\bm{{H}}^{t}_{R}\},t=1,2,...,T$. We design modality-specific decoders for each modality’s features. For unimodal features at different time steps, the decoders are weight-sharing; specifically, all optical features $\bm{{H}}^{t}_{O}$, with added mask tokens and positional embeddings, serve as separate inputs to the optical decoders to reconstruct the original images $\bm{{I}}_{O}^{t}$. SAR modality data are also employed in the same procedure. We compute the mean squared error (MSE) between the reconstructed and original images only for the masked patches of all images [14]. Then, we sum these losses together to form the final loss.
@@ 3.1.5 Decoder
10:         ${\mathbf{H}}^{t+1}_{R}\leftarrow CA(\text{query}=f([\mathbf{F}^{t+1}_{O},\mathbf{F}^{t}_{R}],[\text{key, value}]=\mathbf{F}^{t+1}_{R})$
@@
13:Note: $CA$ denotes the cross-attention layer and $f$ represents linear layers used to fuse features.
3.2 Crop Strategies for Spatial Attribute Mining
Despite the rich geospatial information contained in remote sensing images, the low spatial resolution and the long distances involved result in a low information density of spatial data. Consequently, relying solely on image-level self-supervised learning makes it challenging to learn representations of spatial details. Temporal data capture geographical information of the same location at different times. When inputting them into the network, we can apply different cropping strategies to multi-temporal data. One approach is to use the same crop area for all temporal images (as shown in Figure 2(a)). Alternatively, the multi-temporal images can be partially overlapping spatially (as shown in Figure 2(b)), or the four temporal images can be entirely non-overlapping (as shown in Figure 2(c)).
From the perspective of remote sensing, we posit that, except for changes due to human activities, the geological attributes of the same area in the ecological environment remain constant over time. Therefore, by performing partially overlapping cropping on same-modality images at different times, which means the multi-temporal images input into the network do not completely overlap spatially, we can constrain the network to explore correlations that exist between data in non-overlapping regions. From a computer vision standpoint, CropMAE [39] has demonstrated that using locally correlated cropping in Masked Autoencoder (MAE) tasks can enhance a model’s ability to learn spatial representations.
In Section 4.5.1, we evaluated three cropping methods illustrated in Figure 2. The results indicate that partially overlapping data inputs lead to the best model performance. We believe that using the same region for all temporal images makes the reconstruction task overly simple while employing four entirely non-overlapping images makes it difficult for the model to discover correlations within the data.
@@ 3.3 Progressive pretraining strategy
As depicted in Figure 4 (left), relying solely on a single Masked Image Modeling (MIM) model, the unimodal pretraining procedure lacks interaction across modalities. This approach leads to the model learning multimodal representation implicitly and sub-optimally. Drawing inspiration from models such as GFM [40], SpectralGPT [15], and other MAE-based works [41, 42], we introduce a progressive pretraining strategy for SeaMo. The pretraining process comprises two phases: initially focusing on learning single-time features and subsequently advancing to multi-time features.
@@ 3.3.1 Single-time point multimodal learning
In the initial phase, depicted in Figure 4 (middle), we use data from only a single time point. The first seasonal dataset, which comprises 25% of the total data, is employed for pretraining without temporal information. During this stage, visible tokens from two modalities are concatenated and input into the encoder. As spatial information is crucial for remote sensing interpretation, the focus here is to enable the model to capture both spatial and multimodal features at this singular time point. The foundational understanding of spatial characteristics provides a valuable basis for subsequent phases aimed at developing time-invariant representations.
Figure 4: Three distinct multimodal pretraining strategies. Left: The MIM model is weight-sharing, however, lacks interaction across different modalities. Middle: Data from two modalities are concatenated and then fed into the MIM model. Right: A time series of multimodal images are concatenated (modal-concat) and then processed by the MIM model, followed by a Temporal-Multimodal (TM) block to strengthen representation learning.
@@ 3.3.2 Multi-time flow multimodal learning
In this stage, we initialize the model using pretrained weights from the single-time learning stage. The inputs for this phase include multi-seasonal and multimodal images. Following initial processing by the encoder, Temporal-Multimodal (TM) blocks are integrated to facilitate interactions among seasonal data, as depicted in Figure 4 (right). Within these blocks, multi-seasonal and multimodal tokens are explicitly fused and interact through cross-attention layers in a time flow sequence. TM blocks are essential for enabling the model to acquire time-invariant multimodal representations, significantly enhancing model performance. This functionality has been validated through our ablation studies in Section 4.5. After pretraining, the encoder serves as the backbone for downstream tasks, while the TM blocks and decoders are discarded.
Table 1: Fine-tuning (FT) results on three optical classification benchmarks: EuroSAT, fMoW-S2 and BigEarthNet. On the fMoW-S2 and BEN datasets, we default to using only 10% of the training data to fine-tune the models. * denotes the results obtained by fine-tuning with all training data, rather than only 10% of the data.
@@
Table 2: Linear-probing (LP) results on three optical classification benchmarks: EuroSAT, fMoW-S2 and BigEarthNet (BEN). On the fMoW-S2 and BEN datasets, we default to using only 10% of the training data to fine-tune the models.
@@ 4 Experiments
In this section, we first describe the dataset utilized for pretraining and outline the pretraining setup for our model, SeaMo. Then we report and compare the performance of SeaMo and other models across a range of downstream tasks, including single/multi-label classification, semantic segmentation, and change detection involving various modalities. The results demonstrate SeaMo’s exceptional performance and generalization capabilities. Additionally, we present ablation studies that explore different model configurations, hyperparameter settings, and choices of pretraining strategies.
@@ 4.1.2 Pretraining implementation
We utilize the ViT-Base model as the backbone for our pretraining process. As detailed in Section 3.3, our approach incorporates a progressive pretraining strategy, which enhances data utilization and facilitates the learning of representations. Initially, for the single-time point multimodal learning stage, we use only the first season’s data, which represents 25% of the total dataset, and train without the Temporal-Multimodal (TM) block for 20 epochs. In the subsequent multi-time flow multimodal learning stage, we incorporate the TM block and expand pretraining to include the entire dataset throughout 200 epochs. We utilize the computational resources of 8 NVIDIA GeForce RTX 4090 GPUs for the pretraining of our model. To ensure efficient computing, the images are resized to $128\times 128$ pixels. Consistent with the normalization methods of the SSL4EO-S12 dataset [35], we adopt the same normalization standards. For data augmentation, we apply the partial overlap crop approach mentioned in Section 3.2, each with random horizontal flipping, as implemented in MAE [14]. The batch size is configured at 2048. The training employs the AdamW optimizer [43] with a base learning rate of $1\times 10^{-4}$, which we modulate using a half-cycle cosine decay schedule [44] across 200 epochs.
@@ 4.2 Downstream optical experiments
We evaluate our model SeaMo across multiple optical datasets, focusing on tasks such as single/multi-label classification, change detection, and semantic segmentation. These datasets vary in terms of scale, image size, sample region, and sensor types. The results demonstrate that our model achieves competitive performance relative to various benchmarks. To ensure a fair comparison, we select a range of models for evaluation. The compared methods encompass traditional supervised approaches as well as foundation models that utilize different self-supervised methods and are pretrained on different remote sensing (RS) datasets.
@@ 4.2.1 Single-label classification
We selected the EuroSAT [45] and fMoW-Sentinel[13] datasets for scene classification, with the results displayed in Table 1 and Table 2.
EuroSAT [45] is a dataset comprising 27,000 Sentinel-2 satellite images sourced from 34 European countries. These images are divided into 10 scene classification categories, with each category containing between 2,000 and 3,000 labeled images. Each image in the dataset has a resolution of $64\times 64$ pixels and includes 13 spectral bands. In line with the methodologies used in SatMAE [13] and SpectralGPT [15], we exclude the cirrus band (B10) and utilize the remaining 12 bands for transfer learning. We utilized two NVIDIA GeForce RTX 4090 GPUs for our experiments on the EuroSAT dataset. Images were resized to $128\times 128$ pixels. Data augmentation strategies mirrored those used in [13], including mixup [46] with a coefficient of 0.8, cutmix [47] of 1.0, and a drop path rate [48] of 0.2. The base learning rate was set at $10^{-2}$, with a batch size of 512. Training was conducted over 150 epochs, which included 20 warm-up epochs, utilizing the AdamW optimizer [43]. The model employed soft cross entropy loss as the loss function.
The fMoW-Sentinel dataset presents a more challenging scenario with its 63 categories and global sample strategy, resulting in larger performance discrepancies across models. We utilized only 10% of the training data for fine-tuning. The input images are resized to $128\times 128$ pixels. The base learning rate is set to $2\times 10^{-4}$. The training spanned 80 epochs, which included 20 warm-up epochs. The data augmentation techniques and loss function implementation were aligned with those used for the EuroSAT dataset. Remarkably, our model outperforms all others, including those fine-tuned with the full dataset. Notably, though SatMAE, which is pretrained and then fine-tuned on the fMoW-Sentinel, shows better results than other compared models, SeaMo still surpasses SatMAE [13], demonstrating superior generalization capabilities of our model.

### v2

@@ 3.1.2 Selective masking of patch tokens
A selective masking operation is applied to the token embeddings to differentiate between visible and masked patches. This operation is defined as $[\bm{O}^{t}_{\text{visible}},\bm{O}^{t}_{\text{masked}}]=\mathbb{M}\odot\bm{E}_{O}^{t}\quad\text{and}\quad[\bm{R}^{t}_{\text{visible}},\bm{R}^{t}_{\text{masked}}]=\mathbb{M}\odot\bm{E}_{R}^{t},$ where $\odot$ denotes element-wise multiplication and $\mathbb{M}\in\{0,1\}^{L\times D}$ is a binary mask specifying which tokens are to be masked. Notably, a consistent masking strategy is applied for images from different sensors captured simultaneously, while a random masking strategy is adopted for images acquired at different times. This dual strategy challenges the model to learn more robust and discriminative representations.
3.1.3 Unified multimodal feature encoding
We concatenate the visible tokens from both modalities to form a unified token set for each season $t$. Specifically, given the optical tokens $\bm{O}^{t}_{\text{visible}}\in\mathbb{R}^{L\times D}$ and the SAR tokens $\bm{R}^{t}_{\text{visible}}\in\mathbb{R}^{L\times D},$ we create the combined input: $X^{t}=\text{concat}\left[\bm{O}^{t}_{\text{visible}},\bm{R}^{t}_{\text{visible}}\right]\in\mathbb{R}^{2L\times D},$ where $\text{concat}[\cdot,\cdot]$ denotes concatenation along the token dimension. This combined token set $X^{t}$ is then fed into the encoder $\Phi$, which is composed of standard vision transformer blocks [39]. Within each transformer block, self-attention is computed as follows:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{D}}\right)V,$
where the query, key, and value matrices are obtained by applying learned linear projections on $X^{t}$: $Q=X^{t}W_{Q},K=X^{t}W_{K},V=X^{t}W_{V},$ with $W_{Q}$, $W_{K}$, and $W_{V}$ being the projection matrices.
Through the self-attention mechanism, the encoder effectively captures the inter-token relationships across both modalities and spatial dimensions. As a result, the encoder outputs multimodal features denoted as $\{\bm{F}^{t}_{O},\bm{F}^{t}_{R}\}=\Phi(X^{t}),$ which encapsulate the rich spatial and modality-specific characteristics present in the input data.
@@ 3.1.3 Unified multimodal feature encoding
Figure 3: An illustration of the Temporal-Multimodal (TM) block. In this block, data from each modality not only participate in fusion interactions during the current season but also influence the fusion process in subsequent seasons. For clarity, the symbols in the figure are defined as follows: $k$ denotes the key vector, $q$ denotes the query vector, and $v$ denotes the value vector; $CA$ represents cross-attention; and $f$ indicates the fully connected layer.
3.1.4 Temporal fusion module for multimodal integration
Although the encoder processes seasonal images to yield multimodal features $\{\bm{F}^{t}_{O},\bm{F}^{t}_{R}\}$ for $t=1,2,\ldots,T$, it does not explicitly capture temporal dependencies. To remedy this, we propose the Temporal-Multimodal (TM) fusion block (see Figure 3), inspired by SiamMAE [40]. Unlike traditional cross-attention mechanisms that typically focus on interactions along a single dimension (e.g., between queries and keys from different modalities), the TM block is specifically designed to fuse both temporal and modal information concurrently. In this block, data from each modality not only participate in the fusion process at the current time point but also influence the fusion at subsequent time points.
In practice, the visible tokens of each image serve as queries. Tokens from the other modality of the same season, together with tokens from the previous season, are utilized as keys and values. This joint use of tokens from different modalities and temporal instances enables the TM block to establish cross-attention relationships that integrate both temporal evolution and multimodal interactions, thereby leading to richer and more season-aware representations than conventional cross-attention alone. Algorithm 1 provides the pseudocode for the TM fusion block.
3.1.5 Modality-specific reconstruction modules
The output of the TM fusion block consists of refined features for each modality at every season, denoted as $\{\bm{H}^{t}_{O},\bm{H}^{t}_{R}\}$ for $t=1,2,\ldots,T$. For each modality, we design dedicated reconstruction modules that share weights across seasons. Specifically, the optical features $\bm{H}^{t}_{O}$, augmented with mask tokens and positional embeddings, are input to the optical reconstruction module to regenerate the original images $\bm{I}_{O}^{t}$. The SAR modality follows a similar reconstruction process. The reconstruction loss is computed as the Mean Squared Error (MSE) between the reconstructed and original images, calculated only for the masked patches. The overall loss is obtained by summing the reconstruction losses across all seasons and modalities.
@@ 3.1.5 Modality-specific reconstruction modules
10:         ${\mathbf{H}}^{t+1}_{R}\leftarrow CA(\text{query}=f[\mathbf{F}^{t+1}_{O},\mathbf{F}^{t}_{R}],[\text{key, value}]=\mathbf{F}^{t+1}_{R})$
@@
13:Note: $CA$ denotes the cross-attention layer and $f$ represents the fully connected layers used to fuse features.
3.2 Geospatial region selection strategies
Despite the rich geospatial information contained in remote sensing images, their inherent low spatial resolution and expansive coverage yield a low information density at the image level. As a result, relying solely on image-level self-supervised learning can impede the capture of fine-grained spatial details. In contrast, temporal data provide repeated observations of the same geographical area over different periods, thereby offering an opportunity to enhance spatial representation learning via diverse cropping strategies. For instance, one strategy employs an identical crop region for all temporal images, as illustrated in Figure 2(a). Alternatively, crops may be selected to ensure partial spatial overlap among temporal images, as demonstrated in Figure 2(b). In the most challenging scenario, crops from different temporal images are entirely non-overlapping, as depicted in Figure 2(c).
From a remote sensing perspective, we argue that, aside from changes induced by human activities, the geological attributes of a given area tend to remain consistent over time. By enforcing partially overlapping cropping on same-modality images across different time points, the network is encouraged to learn correlations even in non-overlapping regions. This strategy is conceptually similar to the approach adopted by CropMAE [41], which demonstrated that employing locally correlated cropping in MAE tasks can substantially enhance spatial representation learning. However, while CropMAE primarily focuses on single-image spatial consistency, our method leverages the temporal dimension to further constrain the network and uncover latent cross-temporal correlations. In Section 4.5.1, we evaluate the three cropping methods illustrated in Figure 2. Our experiments indicate that partially overlapping data inputs yield the best performance. We believe that using the identical region for all temporal images oversimplifies the reconstruction task, whereas using entirely non-overlapping crops makes it overly challenging for the network to discover inter-region correlations.
@@ 3.3 Progressive pretraining strategy
As depicted in Figure 4 (left), solely using a single MIM model in a unimodal pretraining framework restricts cross-modal interactions, resulting in implicit and often suboptimal multimodal representations. In contrast, we employ the progressive pretraining strategy, which not only leverages the inherent strengths of spatial and temporal data but also introduces a novel two-phase training scheme that distinctly sets our work apart from previous approaches such as GFM [30], SpectralGPT [2], and other MAE-based methods [42]. By sequentially focusing on single-time point features and then on multi-time flow multimodal learning, our strategy delivers several key advantages:
  • 1.
    Enhanced spatial detail capture: The first phase concentrates on learning fine-grained spatial and multimodal features from a single time point, laying a robust foundation that directly addresses the challenges posed by low spatial resolution and sparse information density in remote sensing images.
  • 2.
    Effective temporal fusion: The second phase explicitly integrates temporal information through TM blocks, enabling the model to fuse cross-temporal data via cross-attention mechanisms. This design is not only novel but also crucial for capturing time-invariant representations, which are often overlooked in conventional MIM frameworks.
  • 3.
    Improved representation quality: By progressively refining the model’s capabilities, from basic spatial understanding to complex temporal integration, our approach achieves a more robust and comprehensive feature representation than methods relying on a single-stage pretraining process.
@@ 3.3.1 Single-time point multimodal learning
As illustrated in Figure 4 (middle), in the initial phase, we focus on learning high-quality spatial and multimodal features from a single temporal snapshot. Specifically, we pretrain the model using only the first seasonal dataset, which accounts for 25% of the total data and intentionally excludes temporal information. During this stage, visible tokens from the two modalities are concatenated and fed directly into the encoder. This approach emphasizes the extraction of fine-grained spatial details and cross-modal correlations, which are critical for remote sensing applications due to the low information density of individual images. By concentrating on a single time point, the model benefits from a simplified training scenario that encourages the learning of robust spatial representations. This foundational stage not only stabilizes the initial learning process but also lays the groundwork for integrating temporal dynamics in the subsequent phase.
Figure 4: Three distinct multimodal pretraining strategies. Left: The MIM model is weight-sharing, however, lacks interaction across different modalities. Middle: Data from two modalities are concatenated and then fed into the MIM model. Right: A time series of multimodal images are concatenated (modal-concat) and then processed by the MIM model, followed by a TM block to strengthen representation learning.
@@ 3.3.2 Multi-time flow multimodal learning
In this stage, we initialize the model using pretrained weights from the single-time learning stage. The inputs for this phase include multi-seasonal and multimodal images. Following initial processing by the encoder, the TM blocks are integrated to facilitate interactions among seasonal data, as depicted in Figure 4 (right). Within these blocks, multi-seasonal and multimodal tokens are explicitly fused and interact through cross-attention layers in a time flow sequence. TM blocks are essential for enabling the model to acquire time-invariant multimodal representations, significantly enhancing model performance. This functionality has been validated through our ablation studies in Section 4.5. After pretraining, the encoder serves as the backbone for downstream tasks, while the TM blocks and decoders are discarded.
Table 1: Fine-Tuning (FT) results on three optical classification benchmarks: EuroSAT, fMoW-S2, and BigEarthNet. On the fMoW-S2 and BEN datasets, we default to using only 10% of the training data to fine-tune the models. * denotes the results obtained by fine-tuning with all training data, rather than only 10% of the data. Best and second-best results are in red and blue, respectively.
@@
Table 2: Linear-Probing (LP) results on three optical classification benchmarks: EuroSAT, fMoW-S2, and BigEarthNet (BEN). On the fMoW-S2 and BEN datasets, we default to using only 10% of the training data to fine-tune the models. Best and second-best results are in red and blue, respectively.
@@ 4 Experiments
In this section, we first describe the dataset utilized for pretraining and outline the pretraining setup for our model, SeaMo. Then we report and compare the performance of SeaMo and other models across a range of downstream tasks, including single/multi-label classification, semantic segmentation, and change detection involving various modalities. The results demonstrate SeaMo’s exceptional performance and generalization capabilities. In addition, we present ablation studies that explore different model configurations, hyperparameter settings, and choices of pretraining strategies.
@@ 4.1.2 Pretraining implementation
We utilize the ViT-Base model as the backbone for our pretraining process. As detailed in Section 3.3, our approach incorporates a progressive pretraining strategy, which enhances data utilization and facilitates the learning of representations. Initially, for the single-time point multimodal learning stage, we use only the first season’s data, which represents 25% of the total dataset, and train without the TM block for 20 epochs. In the subsequent multi-time flow multimodal learning stage, we incorporate the TM block and expand pretraining to include the entire dataset throughout 200 epochs. We utilize the computational resources of 8 NVIDIA GeForce RTX 4090 GPUs for the pretraining of our model. To ensure efficient computing, the images are resized to $128\times 128$ pixels. Consistent with the normalization methods of the SSL4EO-S12 dataset [37], we adopt the same normalization standards. For data augmentation, we apply the partial overlap crop approach mentioned in Section 3.2, each with random horizontal flipping, as implemented in MAE [14]. The batch size is configured at 2048. The training employs the AdamW optimizer with a base learning rate of $1\times 10^{-4}$, which we modulate using a half-cycle cosine decay schedule over 200 epochs.
@@ 4.2 Downstream optical experiments
We evaluate our model, SeaMo, across multiple optical datasets, focusing on tasks such as single/multi-label classification, change detection, and semantic segmentation. These datasets vary in terms of scale, image size, sample region, and sensor types. The results demonstrate that our model achieves competitive performance with respect to various benchmarks. To ensure a fair comparison, we select a range of models for evaluation. The compared methods encompass traditional supervised approaches as well as foundation models that utilize different self-supervised methods and are pretrained on different RS datasets.
@@ 4.2.1 Single-label classification
We selected the EuroSAT [45] and fMoW-Sentinel[13] datasets for scene classification, with results displayed in Tables 1 and 2.
EuroSAT [45] is a dataset comprising 27,000 Sentinel-2 satellite images sourced from 34 European countries. These images are divided into 10 scene classification categories, with each category containing between 2,000 and 3,000 labeled images. Each image in the dataset has a resolution of $64\times 64$ pixels and includes 13 spectral bands. In line with the methodologies used in SatMAE [13] and SpectralGPT [2], we exclude the cirrus band (B10) and utilize the remaining 12 bands for transfer learning. We utilized two NVIDIA GeForce RTX 4090 GPUs for our experiments on the EuroSAT dataset. Images were resized to $128\times 128$ pixels. Data augmentation strategies mirrored those used in [13], including mixup with a coefficient of 0.8, cutmix of 1.0, and a drop path rate of 0.2. The base learning rate has been set to $10^{-2}$, with a batch size of 512. The training has been conducted over 150 epochs, including 20 warm-up epochs, and utilizing the AdamW optimizer. The model employed soft cross-entropy loss as the loss function.
The fMoW-Sentinel dataset presents a more challenging scenario with its 63 categories and global sample strategy, resulting in larger performance discrepancies across models. We used only 10% of the training data for fine-tuning. The input images are resized to $128\times 128$ pixels. The base learning rate was set to $2\times 10^{-4}$. The training spanned 80 epochs, including 20 warm-up epochs. Data augmentation techniques and loss function implementation have been aligned with those used for the EuroSAT dataset. Remarkably, our model outperforms all others, including those fine-tuned with the full dataset. Notably, though SatMAE, which is pretrained and then fine-tuned on the fMoW-Sentinel, shows better results than other compared models, SeaMo still surpasses SatMAE [13], demonstrating superior generalization capabilities of our model.